OpenStudy (anonymous):

Absolute value of quantity 4a +6 = 4a - 10

OpenStudy (watchmath):

You mean like this? \(|4a+6|=4a-10\)

OpenStudy (anonymous):

yes

OpenStudy (watchmath):

Square both sides we have \(16a^2+48a+36=16a^2-80a+100\) \(48a+36=-80a+100\) \(128a=64\) \(a=1/2\) But if we plug in that to the original equation the left side is not the same as the right side. Thus this equation has no solution!.